OpenStudy (anonymous):

Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first. (6161) & (68,69)

OpenStudy (anonymous):

find slope between two points

OpenStudy (anonymous):

Yeah I need someone to solvee it @! :P

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

then use one pt as x1 and y1

OpenStudy (anonymous):

slope = y2-y1/x2-x1 69-61/68-61 8/7

OpenStudy (anonymous):

what about point slope form?

OpenStudy (anonymous):

8/7 is the slope....m now use either points to use point slope formula y - y1 = m (x - x1) i'll use points (61,61) y - 61 = 8/7 (x - 61) y - 61 = 8/7x - 488/7 y = 8/7 x -488/7 + 61 y = 8/7x + 915/7

OpenStudy (anonymous):

sorry i did that wrong....it's y = 8/7x -61/7

OpenStudy (anonymous):

What does the slope of the line represent within the context of your graph?

OpenStudy (anonymous):

8/7x

OpenStudy (anonymous):

and your x intercept is -61/7

OpenStudy (anonymous):

Using the equation that you found, approximately how tall is a person whose arm span is 66 inches

OpenStudy (anonymous):

y = 8/7(66) -61/7 y = 528/7 - 61/7 y = 467/7

OpenStudy (anonymous):

& if my line of best fit are the two points I mentioned in the beggniing of this equatio, then what is the arm span of a 74-inch-tall person

OpenStudy (anonymous):

?

OpenStudy (anonymous):

plug in 74

OpenStudy (anonymous):

to?

OpenStudy (anonymous):

y = 8/7x -61/7 into x

OpenStudy (anonymous):

so multipuly 8/7 by 74 and subtract 621 divided by 7

OpenStudy (anonymous):

y = 8/7(74) - 61/7 y = 592/7 - 61/7 y = 531/7